### v1

@@ 3.1 Pole selection
As already mentioned, the effectiveness of the rational Krylov subspace (7) is strongly related to the quality of the adopted poles $\bm{\xi}_{m-1}=(\xi_{2},\ldots,\xi_{m})^{T}\in\mathbb{C}^{m-1}$ for the problem at hand. Available pole-selection strategies often rely on the properties of the projected matrices – $T_{m}=V_{m}^{T}AV_{m}$ in our case – to compute the next pole $\xi_{m+1}$; see, e.g., [6, 7].
The strategy we propose for the solution of sequences of shifted linear systems does not involve $T_{m}$ at all and is based on the following proposition.
@@
Assume that there exist $j\in\{2,\ldots,m\}$ and $i\in\{1,\ldots,\ell\}$ such that $\xi_{j}=s_{i}$, namely one of the poles employed in the construction of (7) amounts to one of the shifts in (1), and that $b^{(i)}=b$ for every $i=1,\ldots,\ell$. Then,
(17) $x^{(i)}=(A+s_{i}I_{n})^{-1}b\in\mathbf{K}_{m}(A,b,\bm{\xi}_{m-1}).$
Proof 3.3.
An equivalent definition of the rational Krylov subspace (7) is the following
$\mathbf{K}_{m}(A,b,\bm{\xi}_{m-1})=\text{span}\{q_{m-1}(A)p_{m-1}(A)b,\;p_{m-1}\in\mathbb{P}_{m-1}\},$
where $q_{m-1}(x)=\prod_{i=2}^{m}\frac{1}{x+\xi_{i}}$ whereas $\mathbb{P}_{m-1}$ denotes the set of polynomials of degree up to $m-1$; see, e.g., [15]. Since the terms in $q_{m-1}$ commute, we can always select a $\widehat{p}\in\mathbb{P}_{m-2}$ such that
$q_{m-1}(A)\widehat{p}(A)b=(A+\xi_{j}I)^{-1}b=(A+s_{i}I)^{-1}b,$
and the result follows.
Proposition 3.2 implies that if we choose the new pole $\xi_{m+1}$ equal to one of the shifts $s_{i}$ in (1), then $r_{m+1+k}^{(i)}=0$ for any $k\geq 0$. By recalling the underlying low-rank assumption on the solution $X$ to (5), we thus intuitively expect the norm of the residual vectors related to solutions $x^{(j)}$ close to be linearly dependent to $x^{(i)}=(A+s_{i}I)^{-1}b$ to get small as well. This argument suggests the following pole selection strategy: at iteration $m$, once all the residual norms $\|r_{m}^{(j)}\|$ have been computed, we detect the index related to the maximum $\|r_{m}^{(j)}\|$, namely
$\widehat{j}=\operatorname*{arg\,max}_{j\in\Sigma\setminus\Sigma_{C}}\|r_{m}^{(j)}\|,$
and set $\xi_{m+2}=s_{\widehat{j}}$ in line 19 of Algorithm 1.
@@
Also when adopting the pole-selection strategy above we can easily handle the acquisition of new shifts $s_{i}$’s. In particular, once we solve the projected least squares problems related to these new shifts, we just include the corresponding residual norms in the detection of the most suitable new pole.
@@ 3.2 A fully iterative variant
at each iteration. To this end, sparse direct solvers can be employed [29]. However, there are scenarios where an iterative solution may be more suitable. We consider the adoption of a (preconditioned) Krylov method, such as FOM [27] or GMRES [28], for this task, but, in principle, other iterative solvers like multigrid methods can also be employed.
An easy implementation would only see the adoption of the user’s favorite Krylov method in line 6 of Algorithm 1. On the other hand, more sophisticated options would amount, e.g., to the adoption of a space-augmentation approach where we first project (18) onto the already available space $\mathbf{K}_{m}(A,b,\bm{\xi}_{m})$ so as to capitalize on the computational efforts made for its construction. If this space does not contain a sufficiently accurate approximation to the solution to (18), then we can expand this space with some directions belonging to $\mathbf{K}_{k}(A,v_{m})$. Even though this strategy may look appealing, it is unlikely to be successful. Indeed, by keeping an eye on how we selected the pole $\xi_{m+1}$ (cf. section 3.1), the residual $(A+\xi_{m+1}I)V_{m}y-b$ has large norm, namely $\mathbf{K}_{m}(A,b,\bm{\xi}_{m})$ does not contain enough spectral information. It is thus likely that the same happens for the linear system (18).
@@ 4 The case of multiple right-hand sides
For the sake of clarity in the derivation of the method, in the previous sections we considered the right-hand sides $b_{i}$ in (1) to be independent of the shift index $i$. This means that the right-hand side in the matrix equation formulation of the problem, namely equation (5), has rank one. In this section, we show that our procedure does not change significantly by dropping this assumption, given that the right-hand side of (5) remains low-rank. We identify a scenario which is rather common in practice. In particular, we assume that we have a handful of different right-hand sides, say $k$, and, for all of them, we want to solve the shifted linear system in (1) for all the $\ell$ shifts, with $\ell$ much larger than $k$. We are thus interested in problems of the form
@@
where $X=[x^{(1,1)},\ldots,x^{(1,k)},x^{(2,1)},\ldots,x^{(\ell,k)}]\in\mathbb{C}^{n\times k\ell}$, $B=[b^{(1)},\ldots,b^{(k)}]\in\mathbb{R}^{n\times k}$, and $\otimes$ denotes the Kronecker product. We still look for an approximate solution of the form $X_{m}=V_{m}Y_{m}\in\mathbb{C}^{n\times k\ell}$ where now the orthonormal columns of $V_{m}=[\mathcal{V}_{1},\ldots,\mathcal{V}_{m}]\in\mathbb{R}^{n\times mk}$, $\mathcal{V}_{i}\in\mathbb{R}^{n\times k}$ for all $i=1,\ldots,m$, span the block rational Krylov subspace $\mathbf{K}_{m}(A,B,\bm{\xi}_{m-1})=\text{Range}([B,(A+\xi_{2}I)^{-1}B,\ldots,\prod_{i=2}^{m}(A+\xi_{i}I)^{-1}B])$. For this space, the block counterpart of (3) holds true. In particular, we can write
(20) $AV_{m}=V_{m}T_{m}+\xi_{m+1}\mathcal{V}_{m+1}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1}-(I-V_{m}V_{m}^{T})A\mathcal{V}_{m+1}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1},$
where $H_{m}\in\mathbb{R}^{mk\times mk}$ and $\bm{h}_{m+1,m}\in\mathbb{R}^{k\times k}$ come from the block orthogonalization step, and $E_{m}=e_{m}\otimes I_{k}$; see, e.g., [8] for further details.
As before, we are going to exploit (20) for imposing a minimal residual condition to compute the matrix $Y_{m}=[\mathcal{Y}^{(1)}_{m},\ldots,\mathcal{Y}_{m}^{(\ell)}]\in\mathbb{C}^{mk\times k\ell}$, $\mathcal{Y}_{m}^{(i)}\in\mathbb{C}^{mk\times k}$ for all $i=1\,\ldots,\ell$. Following the reasoning presented in section 3, we write
$AV_{m}=$ $\,V_{m+1}\begin{bmatrix}T_{m}\\ \bm{h}_{m+1,m}(\xi_{m+1}I_{k}-\bm{t}_{m+1,m+1})E_{m}^{T}H_{m}^{-1}\end{bmatrix}+(I-V_{m+1}V_{m+1}^{T})A\mathcal{V}_{m+1}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1}$
$=$ $\,V_{m+1}\begin{bmatrix}T_{m}\\ \bm{h}_{m+1,m}(\xi_{m+1}I_{k}-\bm{t}_{m+1,m+1})E_{m}^{T}H_{m}^{-1}\end{bmatrix}+\mathcal{\widetilde{V}}_{m+2}\bm{\alpha}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1}$
$=$ $\,[V_{m+1},\mathcal{\widetilde{V}}_{m+2}]\begin{bmatrix}T_{m}\\ \bm{h}_{m+1,m}(\xi_{m+1}I_{k}-\bm{t}_{m+1,m+1})E_{m}^{T}H_{m}^{-1}\\ \bm{\alpha}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1}\end{bmatrix},$
where $\bm{t}_{m+1,m+1}=\mathcal{V}_{m+1}^{T}A\mathcal{V}_{m+1}$ and $\bm{\alpha}\in\mathbb{R}^{k\times k}$ is the triangular factor coming from the thin-QR factorization of $(I-V_{m+1}V_{m+1}^{T})A\mathcal{V}_{m+1}$, namely $(I-V_{m+1}V_{m+1}^{T})A\mathcal{V}_{m+1}=\mathcal{\widetilde{V}}_{m+2}\bm{\alpha}$. We can thus write the residual matrix $R_{m}=AX_{m}+X_{m}(S\otimes I_{k})-B(\mathbf{1}_{\ell}^{T}\otimes I_{k})$ as
$R_{m}=[V_{m+1},\mathcal{\widetilde{V}}_{m+2}]\left(\begin{bmatrix}T_{m}\\ \bm{h}_{m+1,m}(\xi_{m+1}I_{k}-\bm{t}_{m+1,m+1})E_{m}^{T}H_{m}^{-1}\\ \bm{\alpha}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1}\end{bmatrix}Y_{m}+\begin{bmatrix}I_{mk}\\ 0\\ \end{bmatrix}Y_{m}(S\otimes I_{k})-E_{1}\bm{\beta}(\mathbf{1}_{\ell}^{T}\otimes I_{k})\right),$
where $E_{1}=e_{1}\otimes I_{k}$ and $\bm{\beta}\in\mathbb{R}^{k\times k}$ is such that $B=\mathcal{V}_{1}\bm{\beta}$. Thanks to the orthogonality of $[V_{m+1},\mathcal{\widetilde{V}}_{m+2}]$ we can compute $Y_{m}=[\mathcal{Y}_{m}^{(1)},\ldots,\mathcal{Y}_{m}^{(\ell)}]$ block-wise by solving
$\mathcal{Y}_{m}^{(i)}=\operatorname*{arg\,min}_{\mathcal{Y}\in\mathbb{C}^{mk\times k}}\left\|\begin{bmatrix}T_{m}+s_{i}I_{mk}\\ \bm{h}_{m+1,m}(\xi_{m+1}I_{k}-\bm{t}_{m+1,m+1})E_{m}^{T}H_{m}^{-1}\\ \bm{\alpha}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1}\end{bmatrix}\mathcal{Y}-E_{1}\bm{\beta}\right\|_{F},\quad i=1,\ldots,\ell.$
@@
$[Q_{m}^{(i)},P_{m}^{(i)}]\begin{bmatrix}G_{m}^{(i)}\\ 0\\ \end{bmatrix}=\begin{bmatrix}T_{m}+s_{i}I_{mk}\\ \bm{h}_{m+1,m}(\xi_{m+1}I_{k}-\bm{t}_{m+1,m+1})E_{m}^{T}H_{m}^{-1}\\ \bm{\alpha}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1}\end{bmatrix},$
with $Q_{m}^{(i)}\in\mathbb{C}^{(m+2)k\times mk}$, $P_{m}^{(i)}\in\mathbb{C}^{(m+2)k\times 2k}$, and $G_{m}^{(i)}\in\mathbb{C}^{mk\times mk}$, then
@@
$\|\mathcal{R}_{m}^{(i)}\|_{F}=\left\|\begin{bmatrix}T_{m}+s_{i}I_{mk}\\ \bm{h}_{m+1,m}(\xi_{m+1}I_{k}-\bm{t}_{m+1,m+1})E_{m}^{T}H_{m}^{-1}\\ \bm{\alpha}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1}\end{bmatrix}\mathcal{Y}_{m}^{(i)}-E_{1}\bm{\beta}\right\|_{F}=\|(P_{m}^{(i)})^{*}E_{1}\bm{\beta}\|_{F}.$
It is easy to show that $R_{m}=[\mathcal{R}_{m}^{(1)},\ldots,\mathcal{R}_{m}^{(\ell)}]$ so that $\|R_{m}\|_{F}^{2}=\sum_{i=1}^{\ell}\|\mathcal{R}_{m}^{(i)}\|_{F}^{2}$, and we can thus apply the same pole-selection strategy presented in section 3.1. In particular, we select the next pole $\xi_{m+2}$ for the rational Krylov basis construction as the shift $s_{i}$ related to $\max_{i}\|\mathcal{R}_{m}^{(i)}\|_{F}$. Moreover, we can stop solving for $\mathcal{Y}_{m}^{(i)}$ as soon as we detect a sufficiently small $\|\mathcal{R}_{m}^{(i)}\|_{F}$.
@@
The strategy presented above can also be adopted for problems (1) where we have $\ell$ right-hand sides $b^{(i)}$ that change with the shift $s_{i}$. In this case we have to assume that the matrix $B=[b^{(1)},\ldots,b^{(\ell)}]$ admits a low-rank representation of the form $B=B_{1}B_{2}^{T}$ with $B_{1}\in\mathbb{R}^{n\times k}$, $B_{2}\in\mathbb{R}^{\ell\times k}$ for small $k$. If this is the case, the construction of the block rational Krylov subspace $\mathbf{K}_{m}(A,B_{1},\bm{\xi}_{m-1})=\text{Range}(V_{m})$ can be carried out. On the other hand, the presence of the matrix $B_{2}$ must be considered in the computation of the $Y_{m}$ factor of the approximate solution $X_{m}=V_{m}Y_{m}$. In particular, $Y_{m}=[y_{m}^{(1)},\ldots,y_{m}^{(\ell)}]\in\mathbb{C}^{mk\times\ell}$ can be computed column-wise by solving
$y_{m}^{(i)}=\operatorname*{arg\,min}_{y\in\mathbb{C}^{mk}}\left\|\begin{bmatrix}T_{m}+s_{i}I_{mk}\\ \bm{h}_{m+1,m}(\xi_{m+1}I_{k}-\bm{t}_{m+1,m+1})E_{m}^{T}H_{m}^{-1}\\ \bm{\alpha}\bm{h}_{m+1,m}E_{m}^{T}H_{m}^{-1}\end{bmatrix}y-E_{1}\bm{\beta}B_{2}^{T}e_{i}\right\|.$

### v2

@@ 3.1 Pole selection
As already mentioned, the effectiveness of the rational Krylov subspace (7) is strongly related to the quality of the adopted poles $\bm{\xi}_{m}=(\xi_{1},\ldots,\xi_{m})^{T}\in\mathbb{C}^{m}$ for the problem at hand. Available pole-selection strategies often rely on the properties of the projected matrices $V_{m+1}^{*}AV_{m+1}$ to compute the next pole $\xi_{m+1}$; see, e.g., [Poles1, Poles2].
The strategy we propose for the solution of sequences of shifted linear systems does not involve $V_{m+1}^{*}AV_{m+1}$ solely. It is rather based on the following proposition.
@@
Let $s$ be a shift selected to be a pole during the construction of $V_{m+1}$, the rational Krylov basis matrix, then $x=(A+sI)^{-1}b=V_{m+1}\underline{K}_{m}y$, for some $y\in\mathbb{C}^{m}$.
Proof 3.5.
Let $s$ be a given shift. The rational Arnoldi relation holding at iteration $m$ in Algorithm 1 is given as
$AV_{m+1}\underline{K}_{m}=V_{m+1}\begin{bmatrix}I_{m}-K_{m}\Xi_{m}\\ -\xi_{m}k_{m+1,m}e_{j}^{T}\end{bmatrix}$,
where $\Xi_{m}=\text{diag}(\xi_{1},\ldots,\xi_{m})$. We have
(18) $V_{m+1}\underline{K}_{m}=(A+sI)^{-1}V_{m+1}\underline{T}_{m},$
where $\underline{T}_{m}=\begin{bmatrix}I_{m}-K_{m}(\Xi_{m}-sI)\\ -(\xi_{m}-s)k_{m+1,m}e_{j}^{T}\end{bmatrix}$. Therefore, $(A+sI)^{-1}b=V_{m+1}\underline{K}_{m}y$ for some $y\in\mathbb{C}^{m}$ if and only if $e_{1}=\underline{T}_{m}y$. Consider the last selected pole $s=\xi_{m}$. Unless a happy-breakdown occurs, the rank of the left hand side in (18) is $m$. Hence, we have $\underline{T}_{m}=\begin{bmatrix}T_{m}\\ 0\end{bmatrix}$ and $T_{m}=I_{m}-K_{m}(\Xi_{m}-\xi_{m}I)$ is full rank. Therefore,
$(A+\xi_{m}I)^{-1}b$ $=(A+\xi_{m}I)^{-1}V_{m}e_{1},$
$=(A+\xi_{m}I)^{-1}V_{m}T_{m}T_{m}^{-1}e_{1},$
$=(A+\xi_{m}I)^{-1}V_{m+1}\underline{T}_{m}T_{m}^{-1}e_{1},$
$=V_{m+1}\underline{K}_{m}T_{m}^{-1}e_{1}.$
For other poles, $\xi_{j}$, for $j<m$, notice that $\text{Range}(V_{j+1}\underline{K}_{j})\subseteq\text{Range}(V_{m+1}\underline{K}_{m})$. Combining the last two arguments proves the proposition.
Proposition 3.4 implies that if we choose the new pole $\xi_{m+1}$ equal to one of the shifts $s_{i}$ in (1), then $r_{m+1+k}^{(i)}=0$ for any $k\geq 0$. By recalling the underlying low-rank assumption on the solution $X$ to (5), we expect small norms for the residual vectors corresponding to the solutions that are approximately linearly dependent with solutions that have been included in the search space. This argument suggests the following pole selection strategy: at the beginning of the iterative method, set $\xi_{1}$ to some randomly selected shift from the set of available shifts. For $m\geq 1$, in line 18 of Algorithm 1 perform the following steps:
$\widehat{j}=\operatorname*{arg\,max}_{j\in\Sigma\setminus\Sigma_{C}}\|r_{m}^{(j)}\|,\quad\text{and}\quad\xi_{m+1}=s_{\widehat{j}}$.
We would like to mention that our pole-selection strategy resembles that of some greedy approaches used in reduced basis methods in the context of model order reduction techniques; see, e.g., [JelichBaydounVoigtMarburg2021, QuarteroniRozzaManzoni2011, BaydounVoigtJelichMarburg2020]. However, to the best of our knowledge, it has never been proposed before in our setting.
@@
It is easy to handle new shifts with the pole-selection strategy above. In particular, once we solve the projected least squares problems related to these new shifts, we pick the new pole to be the shift corresponding to the largest residual norm from the previous and new sets of shifts.
@@ 3.2 A fully iterative variant
at each iteration. To this end, sparse direct solvers can be employed [ScottTuma2023]. However, there are scenarios where an iterative solution may be more suitable. We consider the adoption of a (preconditioned) Krylov method, such as FOM [Saad2003] or GMRES [Schultz1986], for this task, but, in principle, other iterative solvers like multigrid methods can also be employed.
An easy implementation would only see the adoption of the user’s favorite Krylov method in line 6 of Algorithm 1. Yet, more sophisticated options would amount, e.g., to the adoption of a space-augmentation approach where we first project (19) onto the already available space $\mathbf{K}_{m+1}(A,b,\bm{\xi}_{m})$ so as to capitalize on the computational efforts made for its construction. If this space does not contain a sufficiently accurate approximation to the solution to (19), then we can expand this space with some directions belonging to $\mathbf{K}_{k}(A,v_{m})$. Even though this strategy may look appealing, it is unlikely to be successful. Indeed, by keeping an eye on how we selected the pole $\xi_{m+1}$ (cf. section 3.1), the residual $(A+\xi_{m+1}I)V_{m}y-b$ has large norm since $\|(A+\xi_{m+1}I)V_{m}y-b\|\geq\|(A+\xi_{m+1}I)V_{m+1}\underline{K}_{m}y-b\|$. It is thus likely that the same happens for the linear system (19).
@@ 4 The case of multiple right-hand sides
For the sake of clarity in the derivation of the method, in the previous sections we considered the right-hand sides $b_{i}$ in (1) to be the same. This means that the right-hand side in the matrix equation formulation of the problem, namely equation (5), has rank one. In this section, we show that our procedure does not change significantly by dropping this assumption, given that the right-hand side of (5) remains low-rank. We identify a scenario which is rather common in practice. In particular, we assume that we have a handful of different right-hand sides, say $k$, and, for all of them, we want to solve the shifted linear system in (1) for all the $\ell$ shifts, with $\ell$ much larger than $k$. We are thus interested in problems of the form
@@
where $X=[x^{(1,1)},\ldots,x^{(1,k)},x^{(2,1)},\ldots,x^{(\ell,k)}]\in\mathbb{C}^{n\times k\ell}$, $B=[b^{(1)},\ldots,b^{(k)}]\in\mathbb{R}^{n\times k}$, and $\otimes$ denotes the Kronecker product. We still look for an approximate solution of the form $X_{m}=V_{m+1}\underline{K}_{m}Y_{m}\in\mathbb{C}^{n\times k\ell}$ where now the orthonormal columns of $V_{m+1}=[\mathcal{V}_{1},\ldots,\mathcal{V}_{m+1}]\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{C}}^{n\times(m+1)k}$, $\mathcal{V}_{i}\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{C}}^{n\times k}$ for all $i=1,\ldots,m+1$, are the first $(m+1)k$ columns of the basis matrix that span the block rational Krylov subspace $\mathbf{K}_{m+1}(A,B,\bm{\xi}_{m})=\text{Range}([B,(A+\xi_{1}I)^{-1}B,\ldots,\prod_{i=1}^{m}(A+\xi_{i}I)^{-1}B])$. For this space, the block counterpart of (8) holds true. In particular, we can write
(21) $AV_{m+1}\underline{K}_{m}=V_{m+1}\underline{H}_{m},$
where $\underline{K}_{m},\underline{H}_{m}\in{\mathbb{C}}^{(m+1)k\times mk}$ are block Hessenberg matrices that result from the block orthogonalization step; see, e.g., [CasR24] for further details. As before, we are going to exploit (21) for imposing a minimal residual condition to compute the matrix $Y_{m}=[\mathcal{Y}^{(1)}_{m},\ldots,\mathcal{Y}_{m}^{(\ell)}]\in\mathbb{C}^{mk\times k\ell}$, $\mathcal{Y}_{m}^{(i)}\in\mathbb{C}^{mk\times k}$ for all $i=1\,\ldots$$\ell$. Following similar steps as in section 3, we have
$R_{m}=V_{m+1}\left(\underline{H}_{m}Y_{m}+\underline{K}_{m}Y_{m}(S\otimes I_{k})-E_{1}\bm{\beta}(\mathbf{1}_{\ell}^{T}\otimes I_{k})\right),$
where $E_{1}=e_{1}\otimes I_{k}$ and $\bm{\beta}\in\mathbb{R}^{k\times k}$ is such that $B=\mathcal{V}_{1}\bm{\beta}$. Since $V_{m+1}$ is unitary, we can compute $Y_{m}=[\mathcal{Y}_{m}^{(1)},\ldots,\mathcal{Y}_{m}^{(\ell)}]$ block-wise by solving
$\mathcal{Y}_{m}^{(i)}=\operatorname*{arg\,min}_{\mathcal{Y}\in\mathbb{C}^{mk\times k}}\left\|(\underline{H}_{m}+s_{i}\underline{K}_{m})\mathcal{Y}-E_{1}\bm{\beta}\right\|_{F},\quad i=1,\ldots,\ell.$
@@
$Q_{m}^{(i)},P_{m}^{(i)}]\begin{bmatrix}G_{m}^{(i)}\\ 0\\ \end{bmatrix}=(\underline{H}_{m}+s_{i}\underline{K}_{m}),$
with $Q_{m}^{(i)}\in\mathbb{C}^{(m+1)k\times mk}$, $P_{m}^{(i)}\in\mathbb{C}^{(m+1)k\times k}$, and $G_{m}^{(i)}\in\mathbb{C}^{mk\times mk}$, then
@@
$\|\mathcal{R}_{m}^{(i)}\|_{F}=\left\|(\underline{H}_{m}+s_{i}\underline{K}_{m})\mathcal{Y}_{m}^{(i)}-E_{1}\bm{\beta}\right\|_{F}=\|(P_{m}^{(i)})^{*}E_{1}\bm{\beta}\|_{F}.$
It is easy to show that $R_{m}=[\mathcal{R}_{m}^{(1)},\ldots,\mathcal{R}_{m}^{(\ell)}]$ so that $\|R_{m}\|_{F}^{2}=\sum_{i=1}^{\ell}\|\mathcal{R}_{m}^{(i)}\|_{F}^{2}$, and we can thus apply the same pole-selection strategy presented in section 3.1. In particular, we select the next pole $\xi_{m+1}$ for the block rational Krylov basis construction as the shift $s_{i}$ related to $\max_{i}\|\mathcal{R}_{m}^{(i)}\|_{F}$. Moreover, we can stop solving for $\mathcal{Y}_{m}^{(i)}$ as soon as we detect a sufficiently small $\|\mathcal{R}_{m}^{(i)}\|_{F}$.
@@
The strategy presented above can also be adopted for problems (1) where we have $\ell$ right-hand sides $b^{(i)}$ that change with the shift $s_{i}$. In this case we have to assume that the matrix $B=[b^{(1)},\ldots,b^{(\ell)}]$ admits a low-rank representation of the form $B=B_{1}B_{2}^{T}$ with $B_{1}\in\mathbb{R}^{n\times k}$, $B_{2}\in\mathbb{R}^{\ell\times k}$ for small $k$. If this is the case, the construction of the block rational Krylov subspace $\mathbf{K}_{m+1}(A,B_{1},\bm{\xi}_{m})$ can be carried out. However, the presence of the matrix $B_{2}$ must be considered in the computation of the $Y_{m}$ factor of the approximate solution $X_{m}=V_{m+1}\underline{K}_{m}Y_{m}$. In particular, $Y_{m}=[y_{m}^{(1)},\ldots,y_{m}^{(\ell)}]\in\mathbb{C}^{mk\times\ell}$ can be computed column-wise by solving
$y_{m}^{(i)}=\operatorname*{arg\,min}_{y\in\mathbb{C}^{mk}}\left\|(\underline{H}_{m}+s_{i}\underline{K}_{m})y-E_{1}\bm{\beta}B_{2}^{T}e_{i}\right\|.$